OpenStudy (anonymous):

Which is an example of the commutative property? A. 21 + (–21) = 0 B. 11 + (–3 + 8) = (11 + –3) + 8 C. –(12 + 7) = –12 + (–7) D. 13 + (–19) = –19 + 13

OpenStudy (anonymous):

@BriannaBear

OpenStudy (anonymous):

Sorry I'm bad at algebra

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@HelpOfTheGods

OpenStudy (igreen):

Commutative Property: \(\sf a + b = b + a\) So what do you think?

OpenStudy (anonymous):

b

OpenStudy (igreen):

No, that's incorrect..that's the Associative Property. \(\sf (a + b) + c = a + (b + c)\)

OpenStudy (anonymous):

c

OpenStudy (igreen):

Don't guess, which one switches the numbers around?

OpenStudy (igreen):

C is the distributive property..it distributes -1 into the parenthesis.

OpenStudy (anonymous):

is it d

OpenStudy (igreen):

Yes, D is correct.

OpenStudy (igreen):

\(\sf 13 + (-19) = -19 + 13\)

OpenStudy (anonymous):

thx